OpenStudy (anonymous):

Jake bought another computer for $3250. This computer depreciates at a rate of 11% every year. Write a model and predict when the computer will be worth $200. My model: 200 = 3250(.89)^x Is this correct? And how do I solve for x (time)?

Nnesha (nnesha):

yep lO_Oks right :-) in order to solve for x first u should divide by 3250 both sides

OpenStudy (anonymous):

@Nnesha I got a really long decimal, so from there I set the decimal equal to (.89)^x, but how do I solve for the exponent? I can't divide by (.89) right?

Nnesha (nnesha):

yep you can't divide by .89 bec of exponent 200/3250 = .0615 \[\huge\rm .0615 = (.89)^x\] now take "ln" both side then you can use power rule power rule \[\large\rm ln x^y = y \ln x\]

OpenStudy (rational):

good question!

OpenStudy (anonymous):

@Nnesha So divide by ln on both sides?

OpenStudy (anonymous):

@Nnesha Oh! I understand now! Thank you very much! And thank you @rational

OpenStudy (rational):

hey sure ?

OpenStudy (rational):

you need to use calculator

OpenStudy (anonymous):

@rational Yes I understand that and I got the correct answer

OpenStudy (rational):

sounds great! whats your final answer ?

OpenStudy (anonymous):

@rational 23.9 years

OpenStudy (rational):

looks perfect!